OpenStudy (anonymous):

Which of the following equations is the result of cross multiplying the given proportion? 3/k+3 = 8/k-2 3k - 2 = 8k + 3 3k - 6 = 8k + 11 3k - 6 = 8k + 24

OpenStudy (netlopes1):

If I understood: \[\dfrac{3}{k+3}=\dfrac{8}{k-2}\rightarrow3(k-2)=8(k+3)\rightarrow 3k-6=8k+24\] ... Would be that?

OpenStudy (anonymous):

I think so:) thank you:))

OpenStudy (netlopes1):

Thank you for your trust